If the Perimeter of a rectangle is 40 and the diagonal is 12 find the area.
Answer by Tanzem(4)

2(l+b)=40
l+b=20 (squaring both the sides)
l2+b2+lb=400....1
Now diagonal= 12
its right angle triangle.
l2+b2= 144......2
substituting 2 in 1 we get
144+lb=400
lb=256 that is area